Finding MSRP. What was the MSRP for a Hummer H1 that sold for $107,272 after an 8% discount?
Answer by checkley71(8403)

107,272-(107,272*.08)=107,272-8581.76=$98,690.24 IS THE REDUCED PRICE AFTER THE 8% DISCOUNT.